# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ A.5 Proof of Proposition 1
Type I error occurs if the model $B\neq A$ generates the text string $\mathbf{Y}_{N}$, but we have
$\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
By triangle inequality, this yields
$h_{N}(B,A)-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq h_{N}(B,B)-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
First, note that by Assumption 2, we have $h_{N}(B,A)-h_{N}(B,B)\geq\epsilon_{1}$. Hence, the type I error occurs only if
$\epsilon_{1}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\leq-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}.$
Equivalently, the type I error occurs only if
$\epsilon_{1}\leq\bigg{|}-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}\bigg{|}+\bigg{|}-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}\bigg{|}.$
So, the type I error only occurs if at least $|-h_{N}(B,A)+\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B,B)+\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}.$
Next, from Theorem 1, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}-h_{N}(B,B)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
Also, from Theorem 2, we have
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq\epsilon_{1}/2\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Therefore, the Type I error is at most
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
$\frac{1}{N}\sum_{n=1}^{N}Z^{B}_{n}<\frac{1}{N}\sum_{n=1}^{N}Z^{A}_{n}.$
By symmetry, the type II error adheres to the same upper bound as the type I error, yielding
$2\exp\bigg{[}-\frac{N(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{1}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N\epsilon_{1}/2}{c_{1}\log(K)}\min\bigg{(}1,\frac{\epsilon_{1}/2}{c_{1}\log(K)}\bigg{)}\bigg{]}.$
A.6 Proof of Proposition 2
Proof.
Type I error occurs if a model $B_{j}\in\mathcal{B}$ generates the text string $\mathbf{Y}_{N}$, but for one model $A_{i}\in\mathcal{A}$ we have
@@ A.6 Proof of Proposition 2
Equivalently, the type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
@@
So, the type I error only occurs if, for at least one of the models $A_{i}\in\mathcal{A}$, we have either $|-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
@@
Next, from Theorem 1, we have
@@
Also, from Theorem 1 we have
@@
A.7 Proof of Lemma 4
@@
Following the same steps as in the proof for Theorem 1, we conclude that $\sum_{n=1}^{N}[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
@@
Next, setting $t=c_{4}N$, we obtain
$\mathbb{P}\bigg{(}\big{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\big{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (13)
@@
First, by Theorem 1 in Reeb and Wolf (2015), we have
$\Big{|}h_{N}(B,B)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\Big{|}\leq\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K).$ (14)
Therefore, we have
@@
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{2D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\log(K)$
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-\frac{1}{N}\sum_{n=1}^{N}\sqrt{D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}}\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}$
$\geq$ $\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}\bigg{(}\sqrt{\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}}-\sqrt{\frac{1}{2N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}}\bigg{)}$
$\geq$ $\frac{1}{2N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{4N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]},$
where the first equality uses the definition of cross-entropy. The first inequality follows from the triangle inequality. The second inequality follows form (14). The third inequality follows from Assumption 3. The fourth inequality follows from QM-AM inequality. The last inequality follows from the difference of squares identity.
@@
$\leq$ $\mathbb{P}\bigg{(}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\leq 2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}-\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}-2c_{4}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq c_{4}\bigg{)}$
$\leq$ $2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]},$
where the first inequality follows from (15). The fourth inequality holds under Assumption 3, which requires $\frac{1}{2N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq 2\log^{2}(K)$ and also under the assumption in the statement of Lemma 4, which requires that $2c_{4}\leq\log^{2}(K)$. Finally, the last inequality follows from (13). This concludes the proof. ∎
A.8 Proof of Proposition 3
@@
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
@@
$\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}.$
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$ (16)
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}.$
From Lemma 4, we know that
$\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (17)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (18)
@@
$2\exp\bigg{[}-\frac{Nc_{4}}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
@@
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t.$
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$

### v4

@@ A.2.1 Proof of Proposition 1
Type I error occurs if a model $B_{j}\in\mathcal{B}$ generates the text string $\mathbf{Y}_{N}$, but for one model $A_{i}\in\mathcal{A}$, we have
@@
Equivalently, type I error occurs only if at least for one of the models $A_{i}\in\mathcal{A}$ we have
@@
So, type I error only occurs if, for at least one of the models $A_{i}\in\mathcal{A}$, we have either $|-h_{N}(B_{j},A_{i})+\frac{1}{N}\sum_{n=1}^{N}Z^{A_{i}}_{n}|>\epsilon_{1}/2$ or $|-h_{N}(B_{j},B_{j})+\frac{1}{N}\sum_{n=1}^{N}Z^{B_{j}}_{n}|>\epsilon_{1}/2$. Hence, we upper bound the type I error as
@@
Next, from Theorem 1-part (a), we have
@@
Also, from Theorem 1 part (a) we have
@@
A.2.2 Proof of Proposition 2
In this section, we first present Lemma 6 that provides an upper bound on the probability that, under Assumption 3, the absolute value of the difference between the cross-entropy of the (unknown) generator model $B$ and the evaluator model $A$ and the entropy of the evaluator model $A$ over string $\mathbf{Y}_{N}$ be less than or equal to a threshold $c_{4}$. In the proof of Proposition 2, we use Lemma 6 as one of the three key inequalities that yield an upper bound on type I error. Note that providing an upper bound on the type II error does not require any assumptions.
Lemma 6.
Under Assumption 3, for any positive constant $c_{4}<\epsilon_{2}$, we have
$\mathbb{P}\bigg{(}|h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})|\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
@@ A.2.2 Proof of Proposition 2
Following the same steps as in the proof for Theorem 1- part (a), we conclude that $\sum_{n=1}^{N}[D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}|\mathbf{Y}_{n-2}]$ is sub-exponential with parameter $S_{N}\in SE(\nu\sqrt{N},\alpha)$, where $\alpha=\nu=-c_{3}\log(\epsilon)$ for a constant $c_{3}>0$. Then, from (10) we have
@@
Next, setting $t=(\epsilon_{2}-c_{4})N$, we obtain
$\mathbb{P}\bigg{(}\big{|}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\sum_{n=1}^{N}\mathbb{E}\Big{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\Big{]}\big{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (14)
@@
First, we have
@@
$\geq$ $\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{|}-h_{N}(A,A)(\mathbf{Y}_{N})$
$\geq$ $\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=3}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}+\epsilon_{2},$
where the first equality uses the definition of cross-entropy. The first and second inequality follow from the triangle inequality. The third inequality follows from Assumption 3.
@@
$\leq$ $\mathbb{P}\bigg{(}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}+\epsilon_{2}\leq c_{4}\bigg{)}$
$=$ $\mathbb{P}\bigg{(}-\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}+\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\geq-c_{4}+\epsilon_{2}\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\bigg{[}D_{KL}\big{(}p_{n}^{B}||p_{n}^{A}\big{)}\bigg{]}\bigg{|}\geq\epsilon_{2}-c_{4}\bigg{)}$
$\leq$ $2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{\epsilon_{2}-c_{4}}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]},$
where the first inequality follows from (15). The one-to-last inequality follows from the property of absolute value. The last inequality follows from (14). This concludes the proof. ∎
Proof of Proposition 2.
@@
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t.$
@@
$\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}.$
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$ (16)
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}-\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\leq t\bigg{)}$
$\leq$ $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}+\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}.$
From Lemma 6, we know that for any positive $c_{4}<\epsilon_{2}$, we have
$\mathbb{P}\bigg{(}\bigg{|}h_{N}(B,A)(\mathbf{Y}_{N})-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\leq c_{4}\bigg{)}\leq 2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (17)
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(B,A)(\mathbf{Y}_{N})\bigg{|}\geq c_{4}-t\bigg{)}\leq 2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$ (18)
@@
$2\exp\bigg{[}-\frac{N(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}-c_{4})}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(c_{4}-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(c_{4}-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
Setting $c_{4}=\epsilon_{2}/2$, yields that the type I error is upper bounded as
$2\exp\bigg{[}-\frac{N(\epsilon_{2}/2)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}/2)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}+2\exp\bigg{[}-\frac{N(\epsilon_{2}/2-t)}{-c_{3}\log(\epsilon)}\min\bigg{(}1,\frac{(\epsilon_{2}/2-t)}{-c_{3}\log(\epsilon)}\bigg{)}\bigg{]}.$
@@
$\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t.$
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\bigg{|}\geq t\bigg{)}\leq 2\exp\bigg{[}-\frac{Nt}{c_{1}\log(K)}\min\bigg{(}1,\frac{t}{c_{1}\log(K)}\bigg{)}\bigg{]},$